The table below lists the numbers of words spoken in a day by each member of randomly selected couples. Complete parts below. Use a 0.05 significant level to test the claim that among couples, males speak fewer words in a day than females.

Female

27000

16870

28783

25177

21677

17178

26813

23831

Male

22665

23117

17592

12991

11902

20167

14957

20671

a. After computing sample mean difference = 5408 and Sd =7097, then the test statistic you use is _____ with degree of freedom is ______
Answer #1

Test statistic,
t = (xbar - mu)/(s/sqrt(n))
t = (5408 - 0)/(7097/sqrt(8))
t = 2.155


degree of freedom = 8 - 1 = 7

This is a paired sample t test